OpenStudy (anonymous):

((a-1)/8) - (3(a-4)/10) = ((2a+4)/5) help please

OpenStudy (anonymous):

solve for a

OpenStudy (anonymous):

\[\frac{ a-1 }{ 8 }-\frac{ 3(a-4) }{ 10 }=\frac{ 2a+4 }{ 5 }\] Your denominators are 8, 10 and 5. What's the smallest number that is divisible by all 3 (lcd)?

OpenStudy (anonymous):

80?

OpenStudy (anonymous):

80's not the smallest, but it will still work :) You need to multiply each of the fractions by 80 so the denominators will cancel when you reduce. \[\frac{ 80(a-1) }{ 8 }-\frac{ (80)(3)(a-4) }{ 10 }=\frac{ 80(2a+4) }{ 5 }\]

OpenStudy (anonymous):

i got 2.215

OpenStudy (anonymous):

as my final answer

OpenStudy (anonymous):

that's not what I got. Can you put up what you did?

OpenStudy (anonymous):

i got -65/23 is that right? @peachpi

OpenStudy (anonymous):

??

OpenStudy (anonymous):

no. the 23 is right, so everything you did with the a is good. check your addition/subtraction with the constants

OpenStudy (anonymous):

11?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

11/23

OpenStudy (anonymous):

can u show me how you got that

OpenStudy (anonymous):